A machine fills boxes with chocolate balls. The weights of the empty boxes are normally distributed with a mean of 150 g and a standard deviation of 10 g. The mean weight of each chocolate ball is 40 g with a standard deviation of 0.8 g. A full box contains 25 chocolate balls.


a. Calculate the mean weight of a full box of chocolate balls

b. Calculate the standard deviation of the weight of a full box of chocolate balls.

c. Calculate the probability that a full box of chocolate balls weighs less than 1140 g.

d. The company also produces boxes of chocolate squares. It was found that 19% of the boxes

weigh less than 800 g and 28% weigh more than 1200 g. Calculate the mean and standard

deviation of the weight of these boxes



1
Expert's answer
2021-06-24T12:32:17-0400

Let "X"and "Y" be independent random variables that are normally distributed (and therefore also jointly so), then their sum is also normally distributed. i.e., if

"X\sim N(\mu_X, \sigma_X^2), Y\sim N(\mu_Y, \sigma_Y^2),"

"W=X+Y,"

Then "W\sim N(\mu_X+\mu_Y, \sigma_X^2+\sigma_Y^2)."

a. Let "X=" the weight of the empty box, "Y_1, Y_2, ...Y_{25}" represent the weights of chocolate balls, "W=X+Y_1+Y_2+...+Y_{25}" represents the weight of of a full box of chocolate balls.

Given "X\sim N(\mu_X, \sigma_X^2), \mu_X=150\ g, \sigma_X=10\ g"


"Y_i\sim N(\mu_{Y_i}, \sigma_{Y_i}^2), \mu_{Y_i}=40\ g, \sigma_{Y_i}=0.8\ g, i=1,2,...,25"

Then

"W\sim N(\mu_W, \sigma_W^2))"

The mean weight of a full box of chocolate balls is


"\mu_W=\mu_X+\mu_{Y_1}+\mu_{Y_2}+...+\mu_{Y_{25}}"

"=150+25(40)=1150 (g)"

b.

The standard deviation of the weight of a full box of chocolate balls is


"\sigma_W=\sqrt{\sigma_W^2}"

"=\sqrt{\sigma_X^2+\sigma_{Y_1}^2+\sigma_{Y_2}^2+...++\sigma_{Y_{25}}^2}"

"=\sqrt{10^2+25(0.8)^2}=2\sqrt{29}\approx10.77033"

c.


"P(W<1140)=P(Z<\dfrac{1140-1150}{2\sqrt{29}})"

"\approx(Z<-0.9284767)\approx0.17658"

d. Let "V" =the weight of the box of chocolate squares: "V\sim N(\mu_V, \sigma_V^2)."


"P(V<800)=P(Z<\dfrac{800-\mu_V}{\sigma_V})=0.19"

"\dfrac{800-\mu_V}{\sigma_V}\approx-0.877896"


"P(V>1200)=1-P(Z\leq\dfrac{1200-\mu_V}{\sigma_V})=0.28"

"\dfrac{1200-\mu_V}{\sigma_V}\approx0.582842"


"\mu_V=800+0.877896\sigma_V"

"1200-800-0.8777896\sigma_V=0.582842\sigma_V"

"\sigma_V\approx273.8541\ g"


"\mu_V\approx1040.4138\ g"